# DriveMoE: Mixture-of-Experts for Vision-Language-Action Model in End-to-End Autonomous Driving

### v1

@@ Abstract
End-to-end autonomous driving (E2E-AD) demands effective processing of multi-view sensory data and robust handling of diverse and complex driving scenarios, particularly rare maneuvers such as aggressive turns. Recent success of Mixture-of-Experts (MoE) architecture in Large Language Models (LLMs) demonstrates that specialization of parameters enables strong scalability. In this work, we propose DriveMoE, a novel MoE-based E2E-AD framework, with a Scene-Specialized Vision MoE and a Skill-Specialized Action MoE. DriveMoE is built upon our $\pi_{0}$ Vision-Language-Action (VLA) baseline (originally from the embodied AI field), called Drive-$\pi_{0}$. Specifically, we add Vision MoE to Drive-$\pi_{0}$ by training a router to select relevant cameras according to the driving context dynamically. This design mirrors human driving cognition, where drivers selectively attend to crucial visual cues rather than exhaustively processing all visual information. In addition, we add Action MoE by training another router to activate specialized expert modules for different driving behaviors. Through explicit behavioral specialization, DriveMoE is able to handle diverse scenarios without suffering from modes averaging like existing models. In Bench2Drive closed-loop evaluation experiments, DriveMoE achieves state-of-the-art (SOTA) performance, demonstrating the effectiveness of combining vision and action MoE in autonomous driving tasks. We will release our code and models of DriveMoE and Drive-$\pi_{0}$.
@@ 1 Introduction
footnotetext: Correspondence author is also affiliated with Shanghai lnnovation Institute. This work was in part supported by by NSFC (62222607) and Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102.
@@
Modern autonomous driving has made significant progress Chitta2023PAMI ; hu2023planning ; jiang2023vad ; wu2023policypretrainingautonomousdriving ; zhu2024flatfusion ; li2023delving with an end-to-end paradigm, which directly maps the raw sensor input into the planning results. This paradigm wu2022trajectoryguided ; Jaeger2023ICCV yields many advantages, such as reduced engineering complexity, mitigation of error propagation, and global objective optimization. Despite the encouraging results achieved on various open-loop self-driving benchmarks song2025don ; wang2025diffad ; jia2024amp ; you2024bench2drive , existing end-to-end models still fail to get satisfactory performance in closed-loop settings shao2023lmdrive ; sima2023drivelm ; chen2024asynchronouslargelanguagemodel . In closed-loop settings, trained driving models can easily encounter out-of-distribution cases zhai2023ADMLP ; zheng2024genad ; jia2021ide ; lu2024activead , requiring stronger generalizability and reasoning abilities.
@@
Figure 1: Comparison of Different Vision and Action Modeling Strategies in VLA-based End-to-End Driving. (a.1) Vanilla visual token encoding sima2023drivelm processes all surround-view images through a vision tower, leading to token redundancy and increased computational cost. (a.2) Query-based token extraction wang2023drivemlm (e.g., Q-former li2023blip ) selects a subset of visual tokens from each image, but loses spatial structure and requires additional pretraining. (a.3) Our proposed Scene-Specialized Vision MoE dynamically selects a subset of cameras—typically frontal and a few context-relevant side/rear views, reducing redundancy. (b.1) Standard action models adopt one policy head to handle all driving scenarios, limiting performance in rare or skill-specific behaviors. (b.2) Our Skill-Specialized Action MoE, built on a flow-matching planner, activates different experts based on driving intent (e.g., lane following, turning, obstacle avoidance), enabling context-aware and behavior-specialized planning.
@@
Vision Language Models(VLM) and Vision Language Action Models(VLA) recently have gain much attention due to their strong generalizability and transferability across domains black2024pi_0 ; intelligence2025pi_ . To enhance generalization and contextual reasoning, recent work Yang2023LLM4DriveAS ; renz2024carllavavisionlanguagemodels ; shao2023lmdrive ; sima2023drivelm has attempted to introduce VLA into autonomous driving. However, existing VLA approaches still face two major limitations. Firstly, existing vision processors of VLA introduce information redundancy and significant computational overhead. As shown in the upper part of Figure 1, there are two distinct strategies for processing multi-view inputs. The first strategy, termed vanilla vision processorshao2023lmdrive ; sima2023drivelm ; chen2024asynchronouslargelanguagemodel , processes all available camera views at each timestep without distinction, resulting in a substantial computational burden and redundant visual representations, thereby limiting efficiency and scalability. The second strategy, termed query-based vision processor, employs learned queries (e.g., Q-former modulesli2023blip ) to extract a compact set of visual tokens guided by semantic context. However, these learned queries typically lead to the loss of precise geometric and positional information and require substantial additional pre-training efforts qi2024gpt4point . Secondly, as shown in the lower part of the Figure 1, current VLA-based frameworks renz2024carllavavisionlanguagemodels ; shao2023lmdrive generally employ a single unified policy network designed to handle the full spectrum of driving behaviors. Such uniform approaches jia2022multi ; jia2023towards ; jia2023hdgt tend to bias model training towards more frequent scenarios, thereby insufficiently addressing rare but critical driving maneuvers, such as emergency braking or aggressive turning. This lack of explicit specialization restricts their effectiveness in dynamically changing and highly context-dependent driving situations. Addressing these two key limitations demands architectural innovations capable of both context-aware dynamic multi-view selection and explicit fine-grained skill specialization.
@@
Meanwhile, Mixture-of-Experts (MoE) architectures cai2025survey ; wan2023efficient have significantly advanced Large Language Models (LLMs) zhu2024llama ; bai2023qwen ; openai2024gpt4technicalreport by partitioning model capacity into multiple expert modules, scaling to larger model sizes without proportional increases in computational demands. Despite their demonstrated success, the extension of MoE principles into the vision and action domains, particularly within autonomous driving, remains largely under-explored. Current end-to-end driving models continue to rely predominantly on unified architectures without explicit dynamic expert selection or specialized behavioral adaptation Yang2023LLM4DriveAS ; renz2024carllavavisionlanguagemodels ; shao2023lmdrive ; sima2023drivelm . This gap motivates exploration into leveraging MoE-based specialization to improve both visual perception and decision-making components in autonomous driving.
@@
To address these challenges, we propose DriveMoE, a novel framework built upon our proposed Drive-$\pi_{0}$, a Vision-Language-Action (VLA) foundation model extended from the embodied AI model $\pi_{0}$ black2024pi_0 . DriveMoE introduces both a Scene-Specialized Vision MoE and a Skill-Specialized Action MoE, specifically designed for end-to-end autonomous driving scenarios. DriveMoE dynamically selects contextually relevant camera views and activates skill-specific experts for specialized planning. The Vision MoE employs a learned router to dynamically prioritize camera views aligned with the immediate driving context, integrating projector layers that fuse these selected views into a cohesive visual representation. This approach mirrors human attentional strategies, allowing efficient processing of only critical visual inputs. Concurrently, the Action MoE leverages another routing mechanism to engage distinct experts within a flow-matching planning architecture lipman2022flow , with each expert dedicated to handling specialized behaviors such as lane following, obstacle avoidance, or aggressive maneuvers. By introducing context-driven dynamic expert selection across both perception and planning modules, DriveMoE ensures efficient resource utilization and robust specialization, significantly improving handling of rare, complex, and long-tail driving behaviors.
The contributions are as follows:
@@
  • We extend the VLA foundation model $\pi_{0}$, originally designed for embodied AI, into the autonomous driving domain, developing Drive-$\pi_{0}$ as a unified framework for visual perception, contextual understanding, and action planning.
@@
2 Related Work
2.1 VLM/VLA in End-to-end Autonomous Driving
The advancement of Large Language Models (LLMs) yang2025trajectory ; fan2025interleave has significantly accelerated the development of Vision-Language Models (VLMs) for autonomous driving. Leveraging powerful generalization, open-set reasoning, and scalability, these models have become influential paradigms for end-to-end driving tasks. Notable examples include DriveGPT-4 xu2024drivegpt4 , LMDrive shao2023lmdrive , and DriveLM sima2023drivelm , which formulate perception and planning tasks as sequences of discrete tokens, enabling better interpretability and facilitating cross-domain knowledge transfer. However, token-based modeling inherently limits the ability to represent continuous control commands and trajectories, which are critical for real-world autonomous driving systems requiring fine-grained control. To address this limitation, the embodied AI community has proposed vision-language-action (VLA) models that represent actions as continuous variables instead of discrete tokens. Methods such as OpenVLA kim2024openvla , Diffusion Policy chi2023diffusion and $\pi_{0}$ black2024pi_0 demonstrate strong performance by modeling continuous action distributions through sequence prediction and global optimization. Nevertheless, these approaches often rely on task-specific policies or instruction-conditioned models, which struggle to generalize across the long-tail distribution of behaviors seen in complex driving environments.
2.2 Mixture-of-Experts in Large Language Models
Sparse Mixture-of-Experts (MoE) architectures have become a mainstream approach for scaling LLMs. By replacing the standard feedforward layers in Transformers with expert modules, models like DeepSeekMoE dai2024deepseekmoe and Mixtral-8x7B dai2024deepseekmoe improve task specialization and representation capacity while maintaining inference efficiency through conditional computation. In robotics, MoE architectures have also been used to address task heterogeneity and long-tailed data distributions. For example, MENTOR huang2024mentor replaces the MLP backbone with MoE layers to enable gradient routing among modular experts, helping mitigate gradient interference in multi-task learning. Despite promising results in language modeling and robot policy learning, the use of MoE in end-to-end autonomous driving remains underexplored.
@@ 3 Method
3.1 Preliminary: Drive-$\pi_{0}$ Baseline
We first establish a strong baseline, Drive-$\pi_{0}$, which builds upon the recently proposed $\pi_{0}$ black2024pi_0 Vision-Language-Action (VLA) framework from embodied AI, and extends it to the domain of end-to-end autonomous driving. As shown in Figure 2, specifically, the input to Drive-$\pi_{0}$ includes: (i) a sequence of surround-view images from onboard multi-camera sensors; (ii) a fixed text prompt (e.g., “Please predict future trajectory”); and (iii) the current vehicle state (e.g., speed, yaw rate, and past trajectory). The network design follows $\pi_{0}$ framework with pre-trained Paligemma VLM beyer2024paligemma as the backbone and a flow-matching-based action module for planned future trajectory generation.
3.2 Motivation: From Drive-$\pi_{0}$ to DriveMoE
With Drive-$\pi_{0}$ as the baseline, we identify two major challenges: (i) adopting VLM to process spatial-temporal surround-view video tokens poses significant challenges to computational resource; (ii) driving performance for rare and difficult scenarios are deficient, even if there is similar data for training. It might be related to the interfere effect of different behaviors, as mentioned in the $\pi_{0}$ paper black2024pi_0 . Inspired by the recent success of Mixture of Experts (MoE) in VLM field jiang2024mixtral ; deepseekai2024deepseekv3technicalreport , we introduce DriveMoE, which extends Drive-$\pi_{0}$ by adding two new Mixture-of-Experts (MoE) modules to tackle the aforementioned challenges: (i) We propose a Scene-Specialized Vision MoE that dynamically selects the most relevant camera views based on the current driving context, effectively reducing redundant visual tokens. (ii) We incorporate a Skill-Specialized Action MoE within a flow-matching transformer to generate more accurate future trajectory distributions tailored to diverse driving skills. Figure 2 illustrates the complete DriveMoE architecture.
@@ 3.3 Scene-Specialized Vision MoE
Typical Vision-Language-Action Models (VLAs) kim2024openvla ; black2024pi_0 usually handle only a single or a few images at a time, whereas autonomous driving must handle multi-view, multi-timestep visual inputs. Concatenating all camera frames into a transformer leads to a visual token bottleneck – an explosion in sequence length that drastically slows training and inference and hampers convergence. Among existing works, shao2023lmdrive ; sima2023drivelm adopts a vanilla vision processor to directly handle all visual tokens, while query-based compression modules (e.g. Q-Former li2023blip ) reduce token count but sacrifice spatial structure, often treating images as a “bag of patches” without fine spatial correspondence qi2024gpt4point .
@@
Formally, we define the image from camera view $v$ at timestep $t$ as $\boldsymbol{I}_{t}^{v}$, where $v\in\{1,2,\dots,N\}$ for $N$ available camera views. In particular, the front-view image at timestep $t$ is denoted by $\boldsymbol{I}^{\text{front}}_{t}$. We introduce a lightweight vision router module $\boldsymbol{R}_{\text{vision}}$, which takes as inputs the front-view embedding $\boldsymbol{e}^{\text{front}}_{t}$ and the future goal waypoint $\boldsymbol{g}_{t}$, computing a probability distribution $\boldsymbol{p}_{t}\in\mathbb{R}^{N}$ across all camera views:
$\boldsymbol{p}_{t}=\text{Softmax}\left(\boldsymbol{R}_{\text{vision}}\left(\boldsymbol{e}^{\text{front}}_{t},\boldsymbol{g}_{t}\right)\right),$ (1)
where each element $p_{t}^{v}$ indicates the selection probability of camera view $v$ at timestep $t$. Notably, this routing happens before the expensive backbone computation, so views not selected can be skipped entirely to save compute. Thus, we obtain the input for VLM:
[…]
We further incorporate learnable positional embeddings (PE) that are unique to each camera viewpoint into their corresponding vision tokens to preserve spatial and positional relation across different camera views. The label for selection of views is annotated by manually designed filters based on future trajectories, bounding box, and maps, detailed in Appendix A. With the annotated binary camera-view selection labels $\boldsymbol{y_{t}}\in\{0,1\}$ , the vision router is trained using the cross-entropy loss:
@@
Figure 3: The Scene-Specialized Vision Mixture-of-Experts.

### v2

@@ Abstract
End-to-end autonomous driving (E2E-AD) demands effective processing of multi-view sensor data and robust handling of diverse and complex driving scenarios, particularly rare maneuvers such as aggressive turns. The recent success of the Mixture-of-Experts (MoE) architectures in Large Language Models (LLMs) demonstrates that expert specialization enables strong scalability. In this work, we propose DriveMoE, a novel MoE-based E2E-AD framework, with a Scene-Specialized Vision MoE and a Skill-Specialized Action MoE. First, we introduce Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$, a Vision-Language-Action (VLA) baseline adapted from Embodied AI for autonomous driving, which serves as the foundation model for DriveMoE. Building on this, we strengthen perception through a carefully designed Vision MoE, where a router adaptively selects context-relevant camera views. This mechanism is inspired by human driving cognition, in which attention is directed to key visual cues rather than to all sensory inputs simultaneously. Beyond perception, we introduce an Action MoE that augments the framework by training a router to activate specialized expert modules tailored to distinct driving behaviors. Within the Action MoE, we implement two distinct styles(Token-level Router and Trajectory-level Router) and extensively explore their applicability in autonomous driving. In Bench2Drive closed-loop evaluations, DriveMoE demonstrates robust performance across diverse driving scenarios, alleviates the mode-averaging effect that limits existing models, and achieves state-of-the-art results with significant improvements over Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$.
footnotetext: This work was supported by Scientific Research Innovation Capability Support Project for Young Faculty (U40) of the Ministry of Education of China (SRICSPYF-ZY2025019). This work was also in part supported by the Science and Technology Commission of Shanghai Municipality (No. 24511103100) and the New Cornerstone Science Foundation through the XPLORER PRIZE.
@@ 1 Introduction
Modern autonomous driving has made significant progress [7, 16, 31, 49, 50, 26, 23, 29, 62, 14] with an end-to-end paradigm, which directly maps the raw sensor input into the planning results. This paradigm [54, 21] offers several advantages, such as reduced engineering complexity, mitigation of error propagation, and global objective optimization. Despite the encouraging results achieved on various open-loop self-driving benchmarks [16, 48, 52], existing end-to-end models still fail to get satisfactory performance in closed-loop settings [45, 47, 5]. In closed-loop settings, trained driving models can easily encounter out-of-distribution cases [60, 63], requiring stronger generalization and reasoning abilities. Vision Language Models(VLM) and Vision Language Action Models(VLA) have recently gained much attention due to their strong generalization and transferability across domains [3, 20, 39, 55, 46, 34, 11]. To enhance generalization and contextual reasoning, recent work [57, 44, 45, 47, 40, 13, 12, 24] has attempted to introduce VLA into autonomous driving.
@@
Figure 1: Comparison of Different Vision and Action Modeling Strategies in VLA-based End-to-End Driving. (a) Vanilla visual token encoding [47] processes all surround-view images through a vision tower, leading to token redundancy and increased computational cost. (b) Query-based token extraction [53] (e.g., Q-Former [35]) selects a subset of visual tokens from each image, but loses spatial structure and requires additional pretraining. (c) Our proposed Scene-Specialized Vision MoE dynamically selects a subset of cameras—typically frontal and a few context-relevant side/rear views, reducing redundancy. (d) Standard action models adopt one policy head to handle all driving scenarios, limiting performance in rare or skill-specific behaviors. (e) Our Skill-Specialized Action MoE, built on a flow-matching planner, activates different experts based on driving intent (e.g., lane following, turning, obstacle avoidance), enabling context-aware and behavior-specialized planning.
@@
However, existing VLA approaches still face two major limitations. Firstly, existing vision processors of VLA introduce information redundancy and significant computational overhead. As shown in the upper part of Figure 1, there are two distinct strategies for processing multi-view inputs. The first strategy, termed vanilla vision processor [45, 47, 5, 18, 19], processes all available camera views at each timestep without distinction, resulting in a substantial computational burden and redundant visual representations, thereby limiting efficiency and scalability. The second strategy, termed query-based vision processor, employs learned queries (e.g., Q-Former modules[35]) to extract a compact set of visual tokens guided by semantic context. However, these learned queries typically lead to the loss of precise geometric and positional information and require substantial additional pre-training efforts [42]. Secondly, as shown in the lower part of the Figure 1, current VLA-based frameworks [44, 45] generally employ a single unified policy network designed to handle the full spectrum of driving behaviors. Such uniform approaches tend to bias model training towards more frequent scenarios, thereby insufficiently addressing rare but critical driving maneuvers, such as emergency braking or aggressive turning. This lack of explicit specialization restricts their effectiveness in dynamically changing and highly context-dependent driving situations. Addressing these two key limitations demands architectural innovations capable of both context-aware dynamic multi-view selection and explicit fine-grained skill specialization.
@@
Meanwhile, Mixture-of-Experts (MoE) architectures [4, 51] have significantly advanced Large Language Models (LLMs) [64, 1, 41] by partitioning model capacity into multiple expert modules, scaling to larger model sizes without proportional increases in computational demands. Despite their demonstrated success, the extension of MoE principles into the vision and action domains, particularly within autonomous driving, remains largely under-explored.
Current end-to-end driving models [28, 43] predominantly rely on predicting trajectory modalities within unified architectures, but lack explicit dynamic expert selection or specialized behavioral adaptation, which limits their scalability. This gap motivates our exploration of MoE-based specialization to enhance both visual perception and decision-making in autonomous driving. Due to limited space, we discuss more details of Related Works in Appendix 5.
@@
To address these challenges, we propose DriveMoE, a novel framework built upon our proposed Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$, a Vision-Language-Action (VLA) foundation model extended from the embodied AI model $\boldsymbol{\pi}_{\boldsymbol{0}}$ [3, 20]. DriveMoE introduces both a Scene-Specialized Vision MoE and a Skill-Specialized Action MoE, carefully designed for end-to-end autonomous driving scenarios. DriveMoE dynamically selects contextually relevant camera views and activates skill-specific experts for specialized planning. The Vision MoE employs a learned router to dynamically prioritize camera views aligned with the immediate driving context, integrating projector layers that fuse these selected views into a cohesive visual representation. This approach mirrors human attentional strategies, allowing efficient processing of only critical visual inputs. Concurrently, the Action MoE leverages another routing mechanism to engage distinct experts within a flow-matching planning architecture [37], with each expert dedicated to handling specialized behaviors such as lane following, obstacle avoidance, or aggressive maneuvers. By introducing context-driven dynamic expert selection across both perception and planning modules, DriveMoE ensures efficient resource utilization and robust specialization, significantly improving handling of rare, complex, and long-tail driving behaviors. The contributions are as follows:
@@
  • We extend the VLA foundation model $\boldsymbol{\pi}_{\boldsymbol{0}}$, originally designed for embodied AI, into the autonomous driving, developing Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ as a unified framework for visual perception, contextual understanding, and action planning.
@@
Figure 3: The Scene-Specialized Vision Mixture-of-Experts.
@@ 2 Method
2.1 Preliminary: Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ Baseline
We first establish a strong baseline, Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$, which builds upon the recently proposed $\boldsymbol{\pi}_{\boldsymbol{0}}$ [3] Vision-Language-Action (VLA) framework from embodied AI, and extends it to the domain of end-to-end autonomous driving. As shown in Figure 2, specifically, the input to Drive-$\pi_{0}$ includes: (i) two consecutive front-view images from onboard multi-camera sensors; (ii) a fixed text prompt (e.g., ”Please predict future trajectory”); and (iii) the current vehicle state (e.g., speed, yaw rate, and past trajectory). The network design follows $\boldsymbol{\pi}_{\boldsymbol{0}}$ framework with pre-trained Paligemma VLM [2] as the backbone and a flow-matching-based action module for planned future trajectory generation.
2.2 Motivation: From Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ to DriveMoE
With Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ as the baseline, we identify two major challenges: (i) adopting VLM to process spatial-temporal surround-view video tokens poses significant challenges to computational resource; (ii) driving performance for rare and difficult scenarios are deficient, even if there is similar data for training. It might be related to the interference effect of different behaviors, as mentioned in the $\boldsymbol{\pi}_{\boldsymbol{0}}$ paper [3]. Inspired by the recent success of Mixture of Experts (MoE) in VLM field [30, 9], we introduce DriveMoE, which extends Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ by adding two Mixture-of-Experts (MoE) modules to tackle the aforementioned challenges: (i) We propose a Scene-Specialized Vision MoE that dynamically selects the most relevant camera views based on the current driving context, effectively reducing redundant visual tokens. (ii) We incorporate a Skill-Specialized Action MoE within a flow-matching transformer to generate more accurate future trajectory distributions tailored to diverse driving skills. Figure 2 illustrates the DriveMoE architecture.
@@ 2.3 Scene-Specialized Vision MoE
Typical Vision-Language-Action Models (VLAs) [33, 3] usually handle only a single or a few images at a time, whereas autonomous driving must handle multi-view, multi-timestep visual inputs. Concatenating all camera frames into a transformer leads to a visual token bottleneck – an explosion in sequence length that drastically slows training and inference and hampers convergence. Among existing works, [45, 47] adopts a vanilla vision processor to directly handle all visual tokens, while query-based compression modules (e.g. Q-Former [35]) reduce token count but sacrifice spatial structure, often treating images as a “bag of patches” without fine spatial correspondence [42].
@@
Formally, we define the image from camera view $v$ at timestep $t$ as $\boldsymbol{I}_{t}^{v}$, where $v\in\{1,2,\dots,N\}$ for $N$ available camera views. In particular, the front-view image at timestep $t$ is denoted by $\boldsymbol{I}^{\text{front}}_{t}$. We introduce a lightweight vision router module $\boldsymbol{R}_{\text{vision}}$, which takes as inputs the front-view embedding $\boldsymbol{e}^{\text{front}}_{t}$ and the future goal waypoint $\boldsymbol{g}_{t}$, computing a probability distribution $\boldsymbol{p}_{t}\in\mathbb{R}^{N}$ across all views:
$\boldsymbol{p}_{t}=\text{S(o}\left(\boldsymbol{R}_{\text{vision}}\left(\boldsymbol{e}^{\text{front}}_{t},\boldsymbol{g}_{t}\right)\right),$ (1)
where each element $p_{t}^{v}$ indicates the selection probability of camera view $v$ at timestep $t$. Notably, this routing happens before the expensive backbone computation, so views not selected can be skipped entirely to save compute. Thus, we obtain the input for VLM: ⟨fixed_view⟩, ⟨fixed_view⟩, ⟨dynamic_view⟩, ⟨dynamic_view⟩, ⟨text⟩, ⟨text⟩.
Figure 4: Token-Level Skill-Specialized Action Mixture-of-Experts.
We further incorporate learnable positional embeddings (PE) that are unique to each camera viewpoint into their corresponding vision tokens to preserve spatial and positional relation across different camera views. The label for selection of views is annotated by manually designed filters based on future trajectories, bounding box, and maps, detailed in Appendix 9. With the annotated binary camera-view selection labels $\boldsymbol{y_{t}}\in\{0,1\}$ , the vision router is trained using the cross-entropy loss: